OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 4.5, 27, 162, 972 ... A. an=4.5(2)n-1 B. an=4.5(6)n-1 C. an=6(2)n-1 D. an=6(4.5)n-1

OpenStudy (***[isuru]***):

just as the earlier one... \[A_{n} = ar^{n-1}\] a = first term = 4.5 r = common ratio = 27/4.5 = 6 so... its \[A_{n} = 4.5 \times 6^{n-1}\] which is the second choice :)